# Ravensburg

### v3

@@ Abstract
GenL is a flexible program that can be used to simulate and/or fit x-ray diffraction data from epitaxial thin films exhibiting Laue oscillations. It utilizes differential evolution within a genetic algorithm for the fitting of data and is based on the kinematic theory of diffraction. Effects of polarization, absorption, the Lorentz factor, as well as instrumental resolution and lattice vibrations are taken into account. Useful parameters that can be extracted after fitting include: atomic interplanar spacings, number of coherently scattering atomic planes, strain profiles along the film thickness, and crystal roughness. The program has been developed in MATLAB and employs a graphical user interface. The deployment strategy is twofold whereby the software can either be obtained in source code form and executed within the MATLAB environment, or as a pre-compiled binary for those who prefer not to run it within MATLAB. Finally, GenL can be easily extended to simulate multilayered film systems, superlattices, and films with atomic steps. The program is released under the GNU General Public Licence.
@@ 1 Introduction and motivation
High crystal quality epitaxial thin films are intriguing model systems. Low defect density and minor crystal misorientation allow for a direct and easy comparison to theoretical predictions [Neugebauer2013]. To experimentally verify the crystal quality of epitaxial thin films, the presence of Laue oscillations in x-ray diffraction (XRD) patterns is regarded as a key indication of high crystallinity [Miller2022]. Laue oscillations are only observable if the coherence length of the epitaxial layer is of the order of the film thickness and the latter is smaller than the longitudinal coherence length of the x-ray photons. The occurrence of Laue oscillations is often used to qualitatively characterize samples, associating their presence with high crystallinity and low defect density [Krauss2008, Shu2020, Song2020, Kamigaki1990, DeTheresa2007], thickness uniformity [Naito2001, Xu2017, Hu2013], smooth and flat interface layering [Abe2002, Xu2017, Miyadera2021, Radu2006], and film homogeneity [Forst2015]. Nevertheless, a quantitative analysis of the Laue oscillations, modeling their angular spacing, intensity decay, and asymmetry around their respective Bragg peak can yield additional structural insights [Fullerton1992, Stierle1993, Komar2017, Miller2022].
@@
For instance, the characteristic period of the Laue oscillations is related to the number of coherently scattering layers, and hence, to the coherent thickness of the layer. However, the thickness determined via the separation of Laue oscillations can deviate from the thickness determined by the spacing of the Kiessig fringes found in reflectivity measurements [Kiessig1931]. A possible origin of this may be the presence of excess material at an interface [Miller2022] with a similar electron density as the film but with a different crystal structure. This region adds to the total film thickness determined via the spacing of the Kiessig fringes but it would not affect the thickness determined via the Laue oscillations as the mismatch in crystal structure terminates the crystal coherence length of the film. The intensity decay of the Laue oscillations can be understood in terms of the roughness of the layer, i.e., the distribution of out-of-plane thickness that contributes to coherent scattering. Moreover, an asymmetry of the Laue oscillation intensity on either side of the Bragg peak is frequently observed [Miller2022] and often associated with out-of-plane strain [Vartanyants2000, Robinson2001, Kastle2002, Komar2017], or other factors [Miller2022]. Numerous powerful open-source program codes exist for simulating and fitting of diffraction patterns including Laue oscillations, e.g., SUPREX [Fullerton1992], GenX [Bjorck2007, Glavic_Bjorck_2022], CADEM [Komar2017], or InteractiveXRDFit [Lichtensteiger2018] alongside with commercially available software, e.g., BedeREFS [Wormington1999] or Leptos [Ulyanenkov2004]. However, the focus of most of them lies mainly on the fitting of superlattice peaks and not on the detailed analysis of the Laue oscillations and their shape. The open-source software SUPREX is designed for superconductor superlattices and written in Fortran and Turbo Pascal [Fullerton1992] with a non-trivial extension to other materials [Komar2017]. GenX is designed to fit x-ray reflectivity patterns [Bjorck2007], being primarily sensitive to electron density profiles along the z-axis. But it can be adjusted to fit high-angle scattering data. Though, reconfiguring the program for these settings is not effortless and detailed analysis of the shape of the oscillations is difficult to implement. CADEM on the other hand, which is written in Matlab, can be used to simulate scattering patterns without a fitting option and without the inclusion of crystal roughness. Finally, InteractiveXRDFit, a Matlab program designed for oxide thin films and heterostructures, utilizes an intermediate approach between the two limits of the kinematic and dynamic x-ray diffraction theories to calculate, but not fit, the diffracted intensity [Lichtensteiger2018]. Besides these, Monte Carlo simulations can be used to reproduce experimental diffraction patterns, however, their application is hindered by their high degree of complexity for modeling realistic sample structures [Komar2017].
@@
Hence, we identified the need for GenL, an open source and easy to adapt Laue oscillation fitting program, which is able to not only simulate but also fit experimental diffraction patterns of epitaxial thin films. In GenL, a representative diffraction pattern, which includes all the aforementioned parameters influencing the shape of the measured oscillations, is simulated and fitted to experimental data. The associated parameters are tuned via a differential evolution algorithm [Storn1997] until a difference minimum of the figure of merit between the data and simulation has been obtained.
@@ 2 Theoretical background
where is the x-ray wavelength and is the angle at which constructive interference occurs between sets of planes associated with the Miller indices . Equivalently, the conditions for Bragg reflections are fulfilled when the photon’s momentum transfer wave vector coincides with the reciprocal crystal lattice point represented by the vector . When operating a diffractometer in a symmetric /2 configuration, i.e., the detector angle and incident angle are varied in such a way that the incident angle is half of the scattering angle 2, the momentum transfer vector becomes one dimensional and perpendicular to the lattice planes (after alignment for miscut and other factors). Its modulus can then be written as:
The length of is varied by changing the scattering angle (angle between source and detector). Assuming the reciprocal lattice vector is oriented parallel to , and referring to this direction as (with being perpendicular to and in the scattering plane), the intensity diffracted from the crystal can be written in the kinematic approximation as:
where corresponds to the total number of coherently scattering lattice planes within the coherence volume of the x-rays, is the position of the atomic plane in the thin film stack (taking in the center of the first layer above the substrate interface), and is the Debye-Waller factor. The element specific Debye-Waller pre-factor includes the contribution of the time-averaged displacement of atoms due to lattice vibrations. The dispersion corrected and -dependent form factor is given by [Wilson1993, Henke1993]:
@@
Several other factors need to be considered [Warren1990, Fullerton1992] if the diffracted intensity is to be compared to the experimentally observed one. The first factor contains the inherent polarization dependence of the scattering of an x-ray with an electron as well as the change in polarization when going through a monochromator. The second factor, , accounts for the absorption of x-rays along the angle dependent path lengths of the incident and diffracted rays. The third factor is the Lorentz factor which takes into account deviations from a perfect instrument and the mosaic nature of the crystal. The experimentally observed intensity is then given by:
@@
where is the electron radius, is the number of elements in the alloy, and is the mole fraction of the element .
Finally, the Lorentz factor captures the angular dependent changes of the intensity due to the combination of monochromator imperfections, beam divergence, finite detection resolution, and mosaicity in the sample. Further details can be found elsewhere [Warren1990]. The combined change in intensity can be corrected via:
@@ 3 GenL
In GenL, a diffraction pattern is calculated based on these equations. However, to model x-ray scattering intensity measured in a non-ideal diffractometer setup of a non-perfect single crystal, additional intensity contributions are included in the calculation. Instrumental peak broadening is accounted for by a Gaussian convolution. Moreover, the background intensity originating from thermal diffuse scattering, fluorescence radiation, Compton scattering, air scattering, the detector dark current, and coherent diffuse scattering from defects, needs to be included. In the close vicinity of a peak, assuming a linear dependence often allows for a reliable modeling of [Birkholz2005]:
@@
where , , and are constants. Both options are available to choose from in GenL. Furthermore, based on the user’s choice, the intensity calculation in GenL can include an additional intensity originating from a single crystalline substrate. Due to the large extension of the substrate compared to the film, the substrate Bragg peak is modeled with a Lorentzian function following:
where is the substrate peak intensity, is the half width at half maximum of the peak and is the expected peak position in . However, for the case that the Bragg peak positions from film and substrate lie close to each other, scattering from film and substrate will interfere, which can cause an asymmetry in the Laue oscillations as well as an intensity shift of the Bragg peak position [Robinson1988]. An accurate calculation of the scattering pattern hence requires to add the scattering amplitudes from film and substrate prior to squaring. It is possible to manually adapt the GenL command line version to simulate a semi-infinite stack of atoms from a substrate, a substrate/film interface, and the atomic stack of the epitaxial layer on top, based on the work presented elsewhere [Als-Nielsen.ch3, Als-Nielsen.ch6]. The calculation of scattering amplitudes from a single crystalline substrate will be part of an upcoming version of the program.
In addition, in GenL, out-of-plane strain in the thin film as well as layer roughness can be taken into account. Strain in a single layer is modeled by displacing the atoms in the calculated one-dimensional stack according to a certain strain profile. The displacement of the atom from its position in an unstrained lattice [Fullerton1992] is given by:
where denotes the out-of-plane distance between the atomic planes for the unstrained case and is a fitting parameter. In the program we consider strain originating from the substrate/film interface () and strain originating from an additional film/capping interface (). The latter is not relevant for sole single layers without capping. The respective induced strain can either be tensile or compressive, resulting in the addition or subtraction of to/from the position in an unstrained lattice, respectively. In GenL it is possible to employ either or both of the strain profiles. The layer roughness is included in GenL employing:
@@
In order to then match the theoretical x-ray scattering pattern calculated in GenL and the measured intensity as best as possible, 13 parameters can be fitted in the program:
@@
GenL is similar to GenX as it also offers a graphical user interface (GUI), both are open-source and can be customized individually for specific applications. Furthermore, both have embedded libraries/interfaces with databases to fetch material and scattering parameters. However, in GenX the layering in a thin film stack is simulated, whereas GenL simulates the atomic ordering. It provides deeper insights into the atomic structures and crystal properties, which GenX fits cannot do. In contrast to GenX, GenL is not applicable to all kinds of layers, e.g., amorphous layers.
An example of such a GenL fit of the scattering pattern of an epitaxial vanadium thin film is displayed in Fig. 1. Based on XRR measurements, the thickness of the V layer is 105 Å. The fit takes the tensile out-of-plane strain due to the negative lattice mismatch between the vanadium and the MgO () substrate into account. The calculated, exponentially decaying strain profile is shown as an inset. The CADEM simulation for the same data set published by Komar et al. [Komar2017] is displayed as a blue line. Also the CADEM simulation includes a strain profile [Komar2017], which is in qualitative agreement with the strain profile presented here. The magnitude of the exponential decay in out-of-plane lattice parameter as a function of position is larger for the fit with GenL, in particular for depths in the sample corresponding to the first two unit cells. The drastic change in is, hence, calculated to occur only close to the substrate/V interface and might include contributions from interface roughness or atomic steps, which may contribute to an asymmetric decay of the oscillations but are not considered here. The over-estimation of the tetragonal unit cell distortion at the interface is therefore considered to be an artifact. Numeric values of fitted strain profiles should be analyzed carefully and may be considered only a qualitative display if important contributions to the scattering intensity, especially at interfaces, have been neglected in the selected model for a certain sample. The CADEM simulation corresponds to a coherent thickness of 84 Å. Based on the GenL fit, the coherent thickness of this V film is 86.3 Å. Komar et al. [Komar2017] attribute this difference to the by XRR determined thickness to a partly oxidized surface with similar electron density but different crystal structure. We see that we can account for the asymmetry of the profile by including a strain profile combined with crystal roughness.

### v4

@@ Abstract
GenL is a flexible program that can be used to simulate and/or fit X-ray reflectivity and X-ray diffraction data from epitaxial thin films exhibiting, for example, Laue oscillations. It utilizes a differential evolution within a genetic algorithm for fitting data and uses a modular approach based on either the kinematic theory of diffraction or the dynamic theory. Effects of polarization, absorption, the Lorentz factor, as well as instrumental resolution and lattice vibrations are taken into account. Useful parameters that can be extracted after fitting include atomic interplanar spacings, number of coherently scattering atomic planes, strain profiles along the film thickness, and crystal roughness. The program has been developed in MATLAB and employs a graphical user interface. The deployment strategy is twofold, whereby the software can either be obtained in source code form and executed within the MATLAB environment, or as a pre-compiled binary for those who prefer not to run it within MATLAB. Finally, GenL can be easily extended to simulate multilayered film systems, superlattices, and films with atomic steps. The program is released under the GNU General Public License.
@@ 1 Introduction and motivation
High crystal quality epitaxial thin films are intriguing model systems. Low defect density and minor crystal misorientation allow for a direct and easy comparison to theoretical predictions [Neugebauer2013]. To experimentally verify the crystal quality of epitaxial thin films, the presence of Laue oscillations in X-ray diffraction (XRD) patterns is regarded as a key indication of high crystallinity [Miller2022]. Laue oscillations are only observable if the coherence length of the epitaxial layer is on the order of the film thickness, and the latter is smaller than the longitudinal coherence length of the X-ray photons. The occurrence of Laue oscillations is often used to qualitatively characterize samples, associating their presence with high crystallinity and low defect density [Krauss2008, Shu2020, Song2020, Kamigaki1990, DeTheresa2007], thickness uniformity [Naito2001, Xu2017, Hu2013], smooth and flat interface layering [Abe2002, Xu2017, Miyadera2021, Radu2006], and film homogeneity [Forst2015]. Nevertheless, a quantitative analysis of the Laue oscillations, modeling their angular spacing, intensity decay, and asymmetry around their respective Bragg peak can yield additional structural insights [Fullerton1992, Stierle1993, Komar2017, Miller2022].
@@
For instance, the characteristic period of the Laue oscillations is related to the number of coherently scattering layers, and hence, to the coherent thickness of the layer. However, the thickness determined via the separation of Laue oscillations can deviate from the thickness determined by the spacing of the Kiessig fringes found in reflectivity measurements [Kiessig1931]. A possible origin of this may be the presence of excess material at an interface [Miller2022] with a similar electron density as the film but with a different crystal structure. This region adds to the total film thickness determined via the spacing of the Kiessig fringes but it would not affect the thickness determined via the Laue oscillations as the mismatch in crystal structure terminates the crystal coherence length of the film. The intensity decay of the Laue oscillations can be understood in terms of the roughness of the layer, i.e., the distribution of out-of-plane thickness that contributes to coherent scattering. Moreover, an asymmetry of the Laue oscillation intensity on either side of the Bragg peak is frequently observed [Miller2022] and often associated with out-of-plane strain [Vartanyants2000, Robinson2001, Kastle2002, Komar2017], or other factors [Miller2022]. Numerous powerful open-source program codes exist for simulating and fitting of diffraction patterns including Laue oscillations, e.g., SUPREX [Fullerton1992], GenX [Bjorck2007, Glavic_Bjorck_2022], CADEM [Komar2017], or InteractiveXRDFit [Lichtensteiger2018] alongside with commercially available software, e.g., BedeREFS [Wormington1999] or Leptos [Ulyanenkov2004]. However, the focus of most of them lies mainly on the fitting of superlattice peaks and not on the detailed analysis of the Laue oscillations and their shape. The open-source software SUPREX is designed for superconductor superlattices and written in Fortran and Turbo Pascal [Fullerton1992] with a non-trivial extension to other materials [Komar2017]. GenX is designed to fit X-ray reflectivity patterns [Bjorck2007], being primarily sensitive to electron density profiles along the z-axis. It is possible to adjust it in order to fit high-angle scattering data. Though, reconfiguring the program for these settings is not effortless and detailed analysis of the shape of the oscillations is difficult to implement. CADEM on the other hand, which is written in Matlab, can be used to simulate scattering patterns without a fitting option and without the inclusion of crystal roughness. Finally, InteractiveXRDFit, a Matlab program designed for oxide thin films and heterostructures, utilizes an intermediate approach between the two limits of the kinematic and dynamic X-ray diffraction theories to calculate, but not fit, the diffracted intensity [Lichtensteiger2018]. Besides these, Monte Carlo simulations can be used to reproduce experimental diffraction patterns, however, their application is hindered by their high degree of complexity for modeling realistic sample structures [Komar2017].
@@
Hence, we identified the need for GenL, an open source and easy to adapt Laue oscillation fitting program, which is able to not only simulate but also fit experimental diffraction patterns of epitaxial thin films. In GenL, a representative diffraction pattern, which includes all the aforementioned parameters influencing the shape of the measured oscillations, is simulated and fitted to experimental data.
The associated parameters are tuned via a differential evolution algorithm [Storn1997] until a minimum difference of the figure of merit between the data and simulation has been obtained.
@@ 2 Theoretical background
where is the X-ray wavelength and is the angle at which constructive interference occurs between sets of planes associated with the Miller indices . Equivalently, the conditions for Bragg reflections are fulfilled when the photon’s momentum transfer wave vector coincides with the reciprocal crystal lattice point represented by the vector . In a symmetric – scan, the goniometer is driven so that the detector moves through while the incident angle remains . Under this condition, the momentum-transfer vector varies only in the direction perpendicular to the lattice planes being probed, assuming proper alignment and correction for miscut and other systematic deviations. Its modulus can then be written as:
2.1 Kinematic approach
Assuming the reciprocal lattice vector is oriented parallel to , and referring to this direction as (with being perpendicular to and in the scattering plane), the intensity diffracted from the crystal can be written in the kinematic approximation as:
where corresponds to the total number of coherently scattering lattice planes within the coherence volume of the X-rays, is the position of the atomic plane in the thin film stack (taking in the center of the first layer above the substrate interface), and is the Debye-Waller factor of layer . This expression is generalized in the code to an arbitrary stack of layers, including a substrate, with spacing between layers, where is an interface index.
The element specific Debye-Waller pre-factor includes the contribution of the time-averaged displacement of atoms due to lattice vibrations. The dispersion corrected and -dependent form factor is given by [Wilson1993, Henke1993]:
@@ 2.1 Kinematic approach
Several other factors need to be considered [Warren1990, Fullerton1992] if the diffracted intensity is to be compared to the experimentally observed one. The first factor includes the inherent polarization dependence of the scattering of an X-ray with an electron, as well as the change in polarization when passing through a monochromator. The second factor accounts for the absorption of X-rays along the angle dependent path lengths of the incident and diffracted rays. The third factor is the Lorentz factor which takes into account deviations from a perfect instrument and the mosaic nature of the crystal. The experimentally observed intensity is then given by:
@@
where is the electron radius, is the number of elements in the alloy, and is the mole fraction of element .
Finally, as an option in the program, the Lorentz factor captures the angular dependent changes of the intensity, due to the combination of monochromator imperfections, beam divergence, finite detection resolution, and mosaicity in the sample and depends on the measurement geometry. Further details can be found elsewhere [Warren1990]. The combined change in intensity can be corrected via:
3 Dynamical approach
The dynamic version of the code is an exact calculation of the X-ray scattering using Holý and Fewster’s formulation [Holy_Fewster]. The essential feature is a matrix propagation using Parratt’s algorithm of the (complex) electron density through the entire stack, including any substrate. In contrast to the algorithm in GenX where slab models are used for each physical layer, here every unit cell is further divided into a number of slabs (typically 100). The algorithm is therefore expected to be much slower than a typical reflectivity calculation, but through a number of optimization steps the run time has been cut by several orders of magnitude compared to a naive implementation, thereby rendering these calculations practical. However, the kinematic approach is still faster and is recommended for initial fits to the data. Furthermore, matrix propagation diverges when large number of slabs are used, but is avoided here by a normalization procedure. Our implementation differs from Holý and Fewster’s in that in their work the electron density is calculated in reciprocal space, whereas we calculate the density in real space. This is to allow for the possibility of strain and roughness, which are easier to conceptualize in real space.
The electron density is calculated [Holy_Fewster] if the type of atoms and their coordinates are known, by overlaying the inverse Fourier transforms of the atomic form factors (including dispersion corrections, absorption and Debye-Waller factors) and dividing by the in-plane area of the unit cell. The average electron density is calculated on a grid of points specified by the user, as well as the the limits of integration of the form factors. Strain is implemented by moving the coordinates prior to the transform. At present, a linear strain profile is included from one or both sides of each layer. Crystal roughness is implemented by a Gaussian weighting procedure of calculations of slabs with a different number of unit cells (truncated at three standard deviations). Monolayers can be implemented as separate unit cells by the user. Finally, a layer of vacuum (or air) is added to both sides of the stack. Once the electron density per unit length has been obtained, the diffraction pattern is calculated by matrix propagation, taking the polarization of the incoming X-ray into account and squaring the reflectivity coefficient at the topmost layer. Thus, reflectivity, absorption, and transmission can be calculated to any precision from an arbitrary stack of layers including the effects of strain and roughness. This includes effects of refraction, Darwin broadening, and total reflection; all effects that are not present in the kinematic approach. We note that the density profile generated by GenL cannot be fed into GenX to calculate the diffraction pattern, owing to the divergences and prohibitively slow calculation. To simulate unpolarized light we perform both the dynamical calculation for s and p polarization and combine them. For the case where the monochromator is located on the incidence side, we obtain:
and a similar expression for when the monochromator is on the detector side. We have omitted the Laue factor for the dynamical case because it diverges at low angles.
Fig. 1 depicts the diffraction pattern using the dynamical approach for a bulk single crystal of GaAs with vacuum on either side. The reflectivity region exactly corresponds to the one calculated from GenX as shown in Fig. 2 and the diffraction pattern exhibits Darwin widths as calculated by Holý and Fewster (inset of Fig. 1). Both s and p polarizations are shown in the figure, in excellent agreement with the calculations in [Ref. [Holy_Fewster]]. Note that the dynamical results deviate from the results from GenX at larger scattering angles, highlighting the potential importance of including diffraction effects when fitting reflectivity.
Moreover, electron density obtained from density functional theory or other codes can be used as direct input into the code and the exact diffraction pattern can be calculated, which can be useful when studying subtle changes to the electron density associated with for example Jahn-Teller distortions or crystal field splitting, and magnetic effects, which are not captured by the use of tabulated form factors.
@@ 4 GenL
In GenL, a diffraction pattern is calculated based on the stated equations. However, to model X-ray scattering intensity measured in a non-ideal diffractometer setup of a non-perfect single crystal, additional intensity contributions are included in the calculation. Instrumental peak broadening is accounted for by a Gaussian convolution. Moreover, the background intensity originating from thermal diffuse scattering, fluorescence radiation, Compton scattering, air scattering, the detector dark current, and coherent diffuse scattering from defects needs to be included. In the close vicinity of a peak, assuming a linear dependence often allows for a reliable modeling of [Birkholz2005]:
@@
where , , and are constants. Both options are available to choose from in GenL. Furthermore, several choices are available for inclusion of the substrate in the intensity calculation in GenL for the kinematic case, . The first choice is to model the substrate Bragg peak with a Lorentzian function following:
where is the substrate peak intensity, is the half width at half maximum of the peak and is the expected peak position in . However, for the case that the Bragg peak positions from film and substrate lie close to each other, scattering from film and substrate will interfere, which can cause an asymmetry in the Laue oscillations as well as an intensity shift of the Bragg peak position [Robinson1988]. An accurate calculation of the scattering pattern hence requires to add the scattering amplitudes from film and substrate prior to squaring. Using the GenL command line version the substrate can be included as any other layer, either in the kinematic or exact dynamic approach. In addition, in GenL, out-of-plane strain in the thin film as well as layer roughness can be taken into account. Strain in a single layer is modeled in the code by displacing the atoms in the calculated one-dimensional stack according to a certain strain profile, which can be either exponential or linear. For the exponential case the displacement of the atom from its position in an unstrained lattice [Fullerton1992] is given by:
where denotes the out-of-plane distance between the atomic planes for the unstrained case and is a fitting parameter. For the linear strain profile, both the slope and the extent into the layer can be fitted. In the program we consider strain originating from the substrate/film interface () and strain originating from an additional film/capping interface (). The latter is not relevant for sole single layers without capping. The respective induced strain can either be tensile or compressive, resulting in the addition or subtraction of to/from the position in an unstrained lattice, respectively. The layer roughness is included in GenL employing:
@@
For a single layer, using the kinematic formulation we have implemented a graphical user interface, whereby the theoretical X-ray scattering pattern calculated in GenL is fitted to the measured intensity using 13 parameters:
@@
GenL is similar to GenX as it also offers a graphical user interface (GUI), both are open-source and can be customized individually for specific applications. Furthermore, both have embedded libraries/interfaces with databases to fetch material and scattering parameters. However, in GenX the layering in a thin film stack is simulated, whereas GenL simulates the atomic ordering. It provides deeper insights into the atomic structures and crystal properties, which GenX fits cannot do. In contrast to GenX, GenL is not applicable to all kinds of layers, e.g., amorphous layers. In cases where information about interatomic spacing and coordination is required, recently developed tools designed to determine pair distribution functions of thin films using laboratory-based X-ray sources can be utilized [one_shot_PDFs].
An example of a GenL fit of the scattering pattern of an epitaxial vanadium thin film is displayed in Fig. 3. Based on XRR measurements, the thickness of the V layer is 105 Å. The fit takes the tensile out-of-plane strain due to the negative lattice mismatch between the vanadium and the MgO () substrate into account. The calculated, exponentially decaying strain profile is shown as an inset. The CADEM simulation for the same data set published by Komar et al. [Komar2017] is displayed as a blue line. Also the CADEM simulation includes a strain profile [Komar2017], which is in qualitative agreement with the strain profile presented here. The magnitude of the exponential decay in out-of-plane lattice parameter as a function of position is larger for the fit with GenL, in particular for depths in the sample corresponding to the first two unit cells. The drastic change in is, hence, calculated to occur only close to the substrate/V interface and might include contributions from interface roughness or atomic steps, which may contribute to an asymmetric decay of the oscillations but are not considered here. The over-estimation of the tetragonal unit cell distortion at the interface is therefore considered to be an artifact. Numeric values of fitted strain profiles should be analyzed carefully and may be considered only a qualitative display if important contributions to the scattering intensity, especially at interfaces, have been neglected in the selected model for a certain sample. The CADEM simulation corresponds to a coherent thickness of 84 Å. Based on the GenL fit, the coherent thickness of this V film is 86.3 Å. Komar et al. [Komar2017] attribute this difference to the by XRR determined thickness to a partly oxidized surface with similar electron density but different crystal structure. We see that we can account for the asymmetry of the profile by including a strain profile combined with crystal roughness.